A bullet moves in an angular direction and its angular velocity w = πt + π in terms of (rad / s) after starting t = 0, its average angular velocity during the time the bullet hits the second round, a few radians per second Is?


1
Expert's answer
2021-02-04T11:42:47-0500

The average angular velocity during the time "\Delta t"


"\omega_{avg}=\frac{1}{\Delta t}\int_0^t\omega(t)dt"


Assume that "t=2(s)"


"\omega_{avg}=\frac{1}{2}\int_0^2(\pi t+\pi)dt=\frac{1}{2}\cdot(\frac{\pi\cdot2^2}{2}+2\pi)=2\pi=6.28(rad/s)" . Answer